OpenStudy (anonymous):

A salsa recipe uses green pepper, onion, and tomato in the extended ratio 2 : 5 : 9. How many cups of onion are needed to make 96 cups of salsa

OpenStudy (anonymous):

96x5 i believe

OpenStudy (whpalmer4):

1 batch makes 2+5+9 = 16 cups of salsa. 96/16 = 6, so you'll need 6*5 = 30 cups of onion to make the salsa.

OpenStudy (anonymous):

thankyou so much!